The thermite reaction is spectacular and exothermic.  Iron(III) oxide, Fe2O3, and metallic aluminum produce molten iron and aluminum oxide in a few seconds, according to the equation(/4):

 Fe2O3(s) + 2 Al(s) → Al2O3(s) + 2 Fe

 (a)  Given the following, calculate ΔH for the thermite reaction.  Show your work.                        2 Al(s) + O2 (g) →Al2O3(s)         ΔH = -1670 kJ

                 2 Fe(s) + O2 (g) → Fe2O3(s)      ΔH = - 822 kJ

           

 

 

 

 

       (b)  How much heat energy will be released when 1000.0 g of iron is produced by this reaction(/4)?




1
Expert's answer
2021-06-15T04:36:01-0400

Gives


"Fe_2o_3(s)+2Al(s)\rightarrow Al_2o_3(s)+2Fe"

"2Al(s)+o_2(s)\rightarrow Al_2o_3(s)" ∆H=-1670KJ

"2Fe+(s)+o_2(g)\rightarrow Fe_2o_3" ∆H=-822kJ


"∆H_{rxn}=1mol\times(-822kJ/mol)+2\times0-(-1670kJ/mol)\times1+2\times0"

"\u2206H_{rxn}=-822+1670=848kJ"